# Modular Compilation for Quantum Chiplet Architectures

### v1

@@ 4.2. Elaboration Stage
(a) A $3$-qubit chiplet with a line topology and connected to a neighboring chiplet by $q_{2}$.
(b) A $3$-qubit subcircuit with $2$ inter-chiplet SWAPs.
[…]
(d) A greedy policy takes the shortest path for routing inter-chiplet gates to their nearest halo qubit.
(e) Qubit routing for the remaining intra-chiplet gates is performed as normal, except with restrictions to ensure inter-chiplet SWAPs are kept static.
(f) The remaining compilation states (basis translation and optimization) are not affected by the additional constraints imposed by inter-chiplet gates.
Figure 3. Showcase of Intra-Chiplet Compilation in the Elaboration Stage.
Following the stratification stage, each chiplet can be compiled almost completely independently from one another. More precisely, the compilation process within each chiplet is largely identical to that of a monolithic quantum architecture. The only exception is the presence of inter-chiplet SWAPs, which must be lowered from the chiplet level to the qubit level. Once this assignment is made, these SWAPs must remain fixed and immutable to subsequent compilation stages. Despite this overhead, restricting the task of compilation to a narrower domain of qubits facilitates both parallelization (as each subcircuit can be processed in parallel) and problem size reduction (as there are less qubits to consider) of the most computationally intensive compilation stages (routing and optimization). To accomplish these tasks, we divide the elaboration stage into two substages: qubit allocation and parallel basis translation and optimization. Figure 3 illustrates this with an example of elaboration for a single chiplet in Figure 3(a) and its corresponding subcircuit in Figure 3(b).
@@ 4.2.1. Qubit Allocation
Like other qubit allocation schemes, we incorporate layout (see Figure 3(c)) and routing (see Figure 3(e)) stages. The presence of inter-chiplet gates necessitates several key changes to the SABRE (sabre_swap) qubit allocation algorithm to achieve correctness and/or maximum parallelism. First, the placement of inter-chiplet gates is a non-parallelizable task (due to cross-chiplet interactions) that strongly influences the layout and routing passes, which could otherwise be fully parallelized across chiplets. To account for this, we introduce a serial stage (shown in Figure 3(d)) between the parallelized layout and routing passes that greedily places inter-chiplet gates on the nearest valid edge between halo qubits. A greedy policy was chosen because inter-chiplet gates have more restrictions on valid placements compared to intra-chiplet gates, while also experiencing substantially higher error rates (smith_2022_chiplets). Once the positions of the inter-chiplet gates are settled, they must remain fixed in place, imposing an additional constraint on qubit routing. To promote this behavior in SABRE, we extended its cost heuristic to weigh the device topology by fidelity. This causes higher error links to assume higher costs, discouraging SABRE from moving the high-error inter-chiplet gates. As a result, SEQC achieves better solutions in the common case. However, there is nothing to prevent SABRE from incorrectly modifying any inter-chiplet SWAPs. Thus, we need to impose hard restrictions to ensure correctness. We establish two disjoint gates sets. One gate set operates strictly within a chiplet. This is the traditional gate set for a monolithic processor. The other gate set is the only one permitted to operate across chiplets, and specifically only implements inter-chiplet SWAPs, i.e., SWAP gates between two halo qubits. By restricting each gate set to its own disjoint part of the overall device topology, we ensure correctness across all cases.
@@ 4.2.2. Parallel Translation and Optimization
The remaining compilation stages of basis translation and circuit optimization can be easily parallelized at the chiplet granularity, as shown in Figure 3(f). Basis translation, in particular, is embarrassingly parallel, as every gate can be translated independently from every other gate. Meanwhile, as discussed in Section 2, optimization algorithms already employ partitioning and parallelization to create tractable subproblems from intractably large circuits, facilitating a smaller problem space, lower memory usage, and faster performance. In modular quantum architectures, we physically motivate partitions through the natural chiplet boundaries, enabling us to reap the same resource and performance benefits of partitioning without compromising on solution quality.
@@ 5. Methodology
To experimentally validate our compiler for modular quantum architectures, we ran a benchmark suite on a wide spectrum of circuit sizes for both SEQC and the strawman baseline. Both SEQC and the baseline compiler were implemented in the Qiskit SDK (qiskit2024) v1.2.41.2.41.2.41.2.4 and Python $3.12.6$. Elaboration time results were taken on a machine with a $128$-core Ampere Altra Max ARM64 processor and $256$ GB of DDR4 RAM. When given a parallelizable workload, SEQC attempts to allocate its tasks evenly among the available physical cores, rounded to the neared whole number. Finally, the simulated annealing technique used in the qubit-to-subcircuit stage uses a starting temperature $T_{0}=200$K, a rate of change of $0.005T$ each round, $\frac{T}{50}$ permutations per round, and a total of $5$ trials per core.
@@ 5.1. Metrics
Non-recurring stratification time. This represents the time taken in the stratification stage, which will be present for our compiler and absent for the baseline. The stratification time represents a one-time overhead to prepare a quantum circuit for execution on a modular architecture. This result can be reused for future executions on that device or any device with an equivalent or more expansive topology.
@@ 5.2. Benchmarks
For our benchmark suite, we leverage a subset of circuits in the Supermarq (supermarq) suite of quantum circuits. The circuits we select are chosen to be representative of practical, real-world applications for quantum computing, while also being feasible to compile for a large number of qubits. In particular, our benchmark suite comprises the following quantum circuits.
@@
[…]
VQE. The variational quantum eigensolver (VQE) (vqe) is another hybrid quantum-classical variational algorithm that is often used in chemistry applications. In our experiments, we choose a $2$-layer ansatz whose parameters are once again randomly generated.
Hamiltonian Simulation. Quantum devices are naturally well-suited to simulating Hamiltonians. The Supermarq benchmark selects the 1D Transverse Field Ising Model (TFIM) system to model, which is relevant for phase transitions in magnetic materials.
@@ 5.3. Simulated Quantum Device Specifications
The topology of the simulated device is composed of a grid lattice of symmetric chiplet modules, where each chiplet features a heavy-hexagon qubit topology (heavy_hex), as shown in Figure 4. The intra-chiplet topology is constructed such that the heavy-hex structure is preserved for the global device topology (ignoring chiplet division). The tessellation requirements of each chiplet restrict valid chiplet sizes to multiples of $10$; however, for our experiments, we limit the size of each chiplet to the minimum of $10$ qubits due to hardware limitations.
Note that for each valid chiplet count, there may exist multiple valid chiplet topologies. In this work, we generate our backends to use the “most-square” topology for a given chiplet count. For example, a $12$-chiplet machine would be generated with a grid of $3\times 4$ chiplets.
@@ 6. Experimental Results
Figure 11. Relative circuit depth of SEQC ormalized to the strawman baseline for devices with 10-qubit chiplets (lower is better).
@@
Relative circuit depth and gate count, shown in Figures 11 and 12, respectively, share similar behavior and analysis. At small circuit sizes, SEQC has higher circuit depth and gate count than the baseline. The gap shrinks as the circuit size increases, although SEQC always tends to perform worse on these metrics on average. This behavior is to be expected, since the baseline directly optimizes for circuit depth and gate count, while SEQC prioritizes other metrics, e.g., reducing inter-chiplet gates.
@@ 7. Related Work
Bandic et al.(bandic2024profiling) also note the benefits of modularity. Their study is motivated by application-specific optimization – they seek techniques that allow a quantum circuit’s structure to influence hardware architecture. In their work, the relationship between circuit structure and mapping efficiency is defined, and this information optimizes the architecture of single- and multi-core quantum processors.
Piveteau and Sutter (circuit_knitting) investigate the benefits of using quasiprobability simulation to allow subcircuits cut from larger quantum circuits to perform their quantum operations independently, simulating halo qubit behavior where the subcircuits connect. When augmented with classical communication between circuits to reduce the simulation overhead incurred by sampling, they demonstrate that resulting circuits reach fixed accuracy with less simulation time. Unlike SEQC that targets quantum compilation, this technique has been developed to partition large quantum circuits into subcircuits that fit on smaller devices, at the cost of a simulation overhead to classically combine them.
@@
Using iterative optimization via the Hungarian algorithm, Hungarian Qubit Assignment (HQA) (escofet2023hungarian) seeks to optimize the qubit-to-core mapping problem. Because the Hungarian algorithm is here extended with the ability to adjust a qubit’s cost matrix according to behavior multiple timeslices ahead, the model can leverage lookahead to optimize further the number of inter-core swaps. However, HQA operates under the assumption that all inter-chiplet communications have the same cost, meaning chiplets must be fully connected. SEQC, on the other hand, is able to work with any arrangement of chiplets and inter-chiplet links (as long as there is a path between any two chiplets), taking into account the distance between pairs of chiplets to generate improved circuits.
@@
Parallelizing SEQC’s recurrent elaboration step was inspired by traditional classical compilation techniques, which has a decades-long history and is now commonplace (Mickunas_1978_ParallelCompilation; Gross_1989_ParallelCompilation). Besides traditional compilation, though, we can also draw parallels between our work and prior research on data movement in the classical space. Data movement is widely understood to be the primary performance bottleneck in the classical architecture space. In the last two decades, work has proliferated on dataflow and spatial architectures, much of which focuses on mapping problems. Consider neural network architectures, for example, which seek to map their physical topologies as closely as possible to problem topologies. Eyeriss (Eyeriss_2016_Chen) approaches the dataflow mapping problem by proposing a Row-Stationary architecture, and then optimizing the mapping of problem to processing element such that elements which are spatially local in the problem are spatially local in hardware during execution. This is not unlike qubit swap optimization, and in fact is performed for many of the same reasons—like inter-chiplet swaps, data movements between more distant classical programming elements in a dataflow architecture are simply more expensive, and minimizing their frequency yields performance improvements.
@@ 8. Conclusion
However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation. Existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set, and existing methods of logical-to-physical qubit placement, routing, unitary synthesis, and quantum circuit optimization, are typically not designed for qubit links of wildly varying levels of duration or fidelity.

### v2

@@ 4.2. Elaboration Stage
(a) An intra-chiplet topology
(b) A $3$-qubit subcircuit post-stratification.
(c) Subcircuit after Intra-Chiplet Layout
(d) Inter-Chiplet SWAP Lowering
(e) Intra-Chiplet Routing
(f) Basis Translation and Optimization
Figure 4. Step-by-step example of intra-chiplet compilation in the Elaboration stage. Gates that cannot be implemented on hardware are shown in red.
Following the stratification stage, each chiplet can be compiled almost completely independently from one another. More precisely, the compilation process within each chiplet is largely identical to that of a monolithic quantum architecture. The only exception is the presence of inter-chiplet SWAPs, which must be lowered from the chiplet level to the qubit level. Once this assignment is made, these SWAPs must remain fixed and immutable to subsequent compilation stages. Despite this overhead, restricting the task of compilation to a narrower domain of qubits facilitates both parallelization (since each subcircuit can be processed in parallel) and problem size reduction (since there are fewer qubits to consider) of the most computationally intensive compilation stages, namely routing and optimization. To accomplish these tasks, we divide the elaboration stage into two substages: qubit allocation and parallel basis translation and optimization. Figure 4 illustrates this with an example of elaboration for a single chiplet in Figure 4(a) and its corresponding subcircuit in Figure 4(b).
@@ 4.2.1. Qubit Allocation
Like other qubit allocation schemes, we incorporate layout (see Figure 4(c)) and routing (see Figure 4(e)) substages. However, the presence of inter-chiplet gates also necessitates several key changes to the SABRE (sabre_swap) qubit allocation algorithm to achieve correctness and/or maximum parallelism.
First, the placement of inter-chiplet gates is a non-parallelizable task (due to cross-chiplet interactions) that strongly influences the layout and routing passes, which could otherwise be fully parallelized across chiplets. To account for this, we introduce a serial stage, as shown in Figure 4(d), between the parallel layout and routing passes that greedily places gates on the nearest valid inter-chiplet edge. A greedy policy was chosen because inter-chiplet gates have more restrictions on valid placements compared to intra-chiplet gates, while also experiencing substantially higher error rates (smith_2022_chiplets).
Once the positions of the inter-chiplet gates are settled, they must remain fixed, imposing an additional constraint on qubit routing. To promote this behavior in SABRE, we extended its cost heuristic to weigh the device topology by fidelity. This causes higher error links to assume higher costs, discouraging SABRE from moving the high-error inter-chiplet gates. As a result, SEQC achieves better solutions in the common case. However, there is nothing to prevent SABRE from incorrectly modifying any inter-chiplet SWAPs. Thus, we need to impose hard restrictions to ensure correctness. We establish two disjoint gates sets. One gate set operates strictly within a chiplet. This is the traditional gate set for a monolithic processor. The other gate set is the only one permitted to operate across chiplets, and specifically only implements inter-chiplet SWAPs. By restricting each gate set to its own disjoint part of the overall device topology, we ensure correctness across all cases.
@@ 4.2.2. Parallel Translation and Optimization
The remaining compilation stages of basis translation and circuit optimization can be easily parallelized at the chiplet granularity, as shown in Figure 4(f). Basis translation, in particular, is embarrassingly parallel, because every gate can be translated independently from every other gate. Meanwhile, as discussed in Section 2, optimization algorithms already employ partitioning and parallelization to create tractable subproblems from intractably large circuits, facilitating a smaller problem space, lower memory usage, and faster performance. In modular quantum architectures, we physically motivate partitions through the natural chiplet boundaries, enabling us to reap the same resource and performance benefits of partitioning without compromising on solution quality.
@@ 5. Methodology
To experimentally validate our compiler for modular quantum architectures, we ran a benchmark suite on a wide spectrum of circuit sizes for both SEQC and the Strawman baseline. Both SEQC and the baseline compiler were implemented in the Qiskit SDK (qiskit2024) v1.2.41.2.41.2.41.2.4 and Python $3.12.6$. Elaboration time results were taken on a machine with a $128$-core Ampere Altra Max ARM64 processor and $256$ GB of DDR4 RAM. When given a parallelizable workload, SEQC attempts to allocate its tasks evenly among the available physical cores, rounded to the neared whole number. Finally, the simulated annealing technique used in the qubit-to-subcircuit stage uses a starting temperature $T_{0}=200$, a rate of change of $0.005T$ each round, $\frac{T}{50}$ permutations per round, and a total of $10$ trials per CPU core.
@@ 5.1. Metrics
Non-recurring stratification time. This represents the time taken in the stratification stage, which will be present for our compiler and absent for the baseline. The stratification time represents a one-time overhead to prepare a quantum circuit for execution on a modular architecture. This result can be reused for future executions on any device with an equivalent or more expansive topology.
@@ 5.2. Benchmarks
(a) BitCode
(b) PhaseCode
(c) GHZ
(d) VQE
(e) Hamiltonian Simulation
Figure 5. Examples of the Benchmark Circuits
For our benchmark suite, we leverage a subset of circuits from the Supermarq (supermarq) suite. The circuits are chosen to be representative of practical, real-world applications for quantum computing, while also being feasible to compile for a large number of qubits. In particular, our benchmark suite comprises the following quantum circuits.
@@
Greenberger–Horne–Zeiling (GHZ). Entanglement is one of the fundamental properties of quantum states and a major source of quantum advantage. The GHZ state represents a maximally entangled $n$-qubit state of $\frac{1}{\sqrt{2}}\left(\ket{|}0>^{\otimes n}+\ket{|}1>^{\otimes n}\right)$. The algorithmic circuit structure is essentially a single chain of CNOT gates, as shown in Figure 5(c) From a compilation perspective, this represents a pure exercise in qubit allocation, because the logical circuit is entirely serial. An $n$-qubit GHZ circuit that fully utilizes a $k$-chiplet device should necessitate $\lfloor\frac{n}{k}\rfloor-1$ logical CNOT operations to extend between chiplets. In turn, this will necessitate at least $2\left(\lfloor\frac{n}{k}\rfloor-1\right)$ inter-chiplet SWAPs to transfer qubit in and out of a chiplet.
VQE. The variational quantum eigensolver (VQE) (vqe) is another hybrid quantum-classical variational algorithm that is often used in chemistry applications. In our experiments, we choose a $2$-layer ansatz whose parameters are again randomly generated, as shown in Figure 5(d). This results in a quantum circuit that is essentially two CNOT chains separated by single-qubit operations. The CNOT chains can be interleaved, which in conjunction with the single-qubit gates, creates many opportunities for parallelism (resembling pipelining in classical computing). Moreover, from a qubit routing perspective, the second CNOT chain forces the compiler to consider how qubit permutations in one circuit layer propagate to future layers.
Hamiltonian Simulation. Quantum devices are naturally well-suited to simulating Hamiltonians. The Supermarq benchmark selects the 1D Transverse Field Ising Model (TFIM) system to model, which is relevant for phase transitions in magnetic materials. The circuit for Hamiltonian Simulation, shown in Figure 5(e), resembles a slightly more complex GHZ circuit, including how it is completely serial. Compared to GHZ, this circuit should test the compiler’s ability to lookahead during routing to detect back-to-back gates on the same two-qubit pair, while also emphasizing parallelism originating from the efficient placement of SWAP gates.
@@ 5.3. Simulated Quantum Device Specifications
The topology of the simulated device is composed of a grid lattice of symmetric chiplet modules, where each chiplet features a heavy-hexagon qubit topology (heavy_hex), as shown in Figure 6. The intra-chiplet topology is constructed such that the heavy-hex structure is preserved for the global device topology (ignoring chiplet division). The tessellation requirements of each chiplet restrict valid chiplet sizes to multiples of $10$; however, for our experiments, we limit the size of each chiplet to the minimum of $10$ qubits due to hardware limitations. Finally, note that for each valid chiplet count, there may exist multiple valid chiplet topologies. In this work, we generate our backends to use the “most-square” topology for a given chiplet count. For example, a $12$-chiplet machine would be generated with a grid of $3\times 4$ chiplets.
@@ 6. Experimental Results
Figure 13. Relative circuit depth of SEQC normalized to the Strawman baseline for devices with 10-qubit chiplets (lower is better).
@@
Relative circuit depth and gate count, shown in Figures 13 and 14, respectively, share similar behavior and analysis. At small circuit sizes, SEQC has higher circuit depth and gate count than the baseline. The gap shrinks as the circuit size increases, although SEQC always tends to perform worse on these metrics on average. This behavior is to be expected, since the baseline directly optimizes for circuit depth and gate count, while SEQC prioritizes other metrics, e.g., reducing inter-chiplet gates.
@@ 7. Related Work
Bandic et al.(bandic2024profiling) also note the benefits of modularity. Their study is motivated by application-specific optimization – they seek techniques that allow a quantum circuit’s structure to influence hardware architecture. They define the relationship between circuit structure and mapping efficiency, utilizing this information to optimize the architecture of single- and multi-core quantum processors.
Piveteau and Sutter (circuit_knitting) investigate the benefits of using quasiprobability simulation to allow subcircuits cut from larger quantum circuits to perform their quantum operations independently, simulating inter-chiplet link behavior where the subcircuits connect. When augmented with classical communication between circuits to reduce the simulation overhead incurred by sampling, they demonstrate that resulting circuits reach fixed accuracy with less simulation time. Unlike SEQC that targets quantum compilation, this technique has been developed to partition large quantum circuits into subcircuits that fit on smaller devices, at the cost of a simulation overhead to classically combine them.
@@
Using iterative optimization via the Hungarian algorithm, Hungarian Qubit Assignment (HQA) (escofet2023hungarian) seeks to optimize the qubit-to-core mapping problem. Because the Hungarian algorithm is here extended with the ability to adjust a qubit’s cost matrix according to behavior multiple timeslices ahead, the model can leverage lookahead to optimize further the number of inter-core swaps. However, HQA operates under the assumption that all inter-chiplet communications have the same cost, meaning chiplets must be fully connected. SEQC, on the other hand, is able to work with any arrangement of chiplets and inter-chiplet links so long there is a path between any two chiplets, taking into account the distance between pairs of chiplets to generate improved circuits.
@@
Parallelizing SEQC’s recurrent elaboration step was inspired by traditional classical compilation techniques, which has a decades-long history and is now commonplace (Mickunas_1978_ParallelCompilation; Gross_1989_ParallelCompilation). Besides traditional compilation, though, we can also draw parallels between our work and prior research on data movement in the classical space. Data movement is widely understood to be the primary performance bottleneck in the classical architecture space. In the last two decades, work has proliferated on dataflow and spatial architectures, much of which focuses on mapping problems. Consider neural network architectures, for example, which seek to map their physical topologies as closely as possible to problem topologies. Eyeriss (Eyeriss_2016_Chen) approaches the dataflow mapping problem by proposing a row-stationary architecture, and then optimizing the mapping of problem to processing element such that elements which are spatially local in the problem are spatially local in hardware during execution. This is not unlike qubit swap optimization, and in fact is performed for many of the same reasons — like inter-chiplet swaps, data movements between more distant classical programming elements in a dataflow architecture are more expensive, and minimizing their frequency yields performance improvements.
@@ 8. Conclusion
However, the complexity of chiplet-based quantum devices, coupled with their growing size, presents an imminent scalability challenge for quantum compilation. Existing qubit allocation methods are often unable to contend with inter-chiplet links, which don’t necessary support a universal basis gate set, and existing methods of logical-to-physical qubit placement, routing, unitary synthesis, and quantum circuit optimization, are typically not designed for qubit links that vary significantly in terms of duration or fidelity.
@@
Acknowledgements
This research was partially supported by NSF awards CNS-1763743 and CCF-2119069, Academic Year Undergraduate Research Award #2023-17704 by Northwestern University, and undergraduate summer research awards by the Robert R. McCormick School of Engineering and Applied Science and the Computer Science Department at Northwestern University. The authors thank Kabir Dubey for helping with earlier versions of this manuscript.
This research was supported in part through the computational resources and staff contributions provided for the Quest high performance computing facility at Northwestern University which is jointly supported by the Office of the Provost, the Office for Research, and Northwestern University Information Technology.
We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.